OpenStudy (anonymous):

I have a few questions about this equation>> y=3sin(6(x+2))+4 I know that the amplitude is 3 How would I find what the period is? What is the horizontal shift? (left or right? and by how much?) How would I find the midline?

OpenStudy (anonymous):

@tkhunny I closed this on accident. I reopened it in the "Open Questions" section if you cannot reply on this one

OpenStudy (tkhunny):

Think on \(y = \sin(x)\) Amplitude: 1 Period : \(2\pi\) Midline: y = 0 Horizontal Shift; None You recognized that \(y = 3\sin(x)\) results in an amplitude of 3. Amplitude: 3 Period : \(2\pi\) Midline: y = 0 Horizontal Shift; None The midline shift is the only other easy part. That 4 is just kind of hanging out there. Once you see it, you are not likely to forget it. Thus \(y = 3\sin(x) + 4\) results in Amplitude: 3 Period : \(2\pi\) Midline: y = 4 Horizontal Shift; None Are we making sense, so far?